if 146.5 g of chromium chlorite decomposes, what mass of oxygen can be produced


1
Expert's answer
2022-02-25T15:31:05-0500

Solution:

chromium(III) chlorate = Cr(ClO3)3


Calculate the moles of Cr(ClO3)3:

The molar mass of Cr(ClO3)3 is 302.35 g/mol

Therefore,

Moles of Cr(ClO3)3 = (146.5 g) × (1 mol / 302.35) = 0.4845 mol


Balanced chemical equation:

2Cr(ClO3)3 → 2CrCl3 + 9O2

According to stoichiometry:

2 mol of Cr(ClO3)3 produces 9 mol of O2

Thus, 0.4845 mol of Cr(ClO3)3 produces:

Moles of O2 = [0.4845 mol Cr(ClO3)3] × [9 mol O2 / 2 mol Cr(ClO3)3] = 2.18025 mol O2


Calculate the mass of O2:

The molar mass of O2 is 32.0 g/mol

Therefore,

Mass of O2 = (2.18025 mol O2) × (32.0 g O2 / 1 mol O2) = 69.768 g O2 = 69.77 g O2

Mass of O2 = 69.77 g


Answer: 69.77 grams of oxygen (O2) can be produced.